# Grassroots Platforms with Atomic Transactions: Social Networks, Cryptocurrencies, and Democratic Federations

### v2

@@ 4.3. The Three Platforms are Grassroots as Specified
Proof.
Consider $P\subset P^{\prime}$ and a configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$. This means that all friendships of members of $P$ in $c$ are with other members in $P$. Hence the transaction in $R(P^{\prime})$ in which a member of $P$ establishes friendship with a member of $P^{\prime}\setminus P$ satisfies the definition of interactivity. ∎
@@
Proof.
Consider $P\subset P^{\prime}$ and a configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$. This means that all coins held by members of $P$ in $c$ are $P$-coins. Hence the transaction in $R(P^{\prime})$ in which a member of $P$ swaps coins with a member of $P^{\prime}\setminus P$ (possibly preceded by transactions in which the two members mint coins, if they have not done so already) satisfies the definition of interactivity. ∎
@@
Proof.
Consider $P\subset P^{\prime}$ and a configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$. This means that the subgraph of $G$ projected by members of $P$ is a connected component of $G$. Hence the Join transaction in $R(P^{\prime})$ in which a member of $P$ forms an edge with a member of $P^{\prime}\setminus P$, the direction of which depends on the order $\succ$ of their identifiers, satisfies the definition of interactivity. ∎
5. Implementation
The notion of implementations among distributed transition systems, as well as their fault-tolerance, has been studied extensively (Abadi and Lamport, 1993; Lamport, 1999; Menezes et al., 1995; Manolios, 2003; Shapiro, 2021; Wilcox et al., 2015). Here, we discuss this notion briefly and informally, and expect follow-on work to do so formally.
@@ 5. Implementation
A challenge in this implementation is that a faulty $p$ may fail to either commit or abort following an accept by $q$, leaving $q$ in limbo, at least in regards to this transaction. We will discuss standard and non-standard solutions to this in a subsequent publication.
@@
$k$-ary transactions. The $k$-ary transaction, in which two communities joined or one community leaves another, calls for a consensus protocol executed by the members of the community. To a first approximation, any permissioned consensus (Byzantine Atomic Broadcast) protocol would do. However, the transactions entail changes in the set of agents that participate in consensual decisions, known as “reconfiguration” in the consensus literature. More generally, the grassroots platform entails multiple dynamically-changing, partially-overlapping, sets of agents engaged in running partially-dependent consensus protocols. This challenge seems to have not been previously addressed in a principled way.
6. Related and Future Work
Atomic transactions have been investigated early in distributed computing, mostly in the context of database systems (Lampson, 1981; Lynch and Merritt, 1993; Lynch et al., 1988). Most research since and until today focuses on their efficient and robust implementation (Bravo and Gotsman, 2019; Chockler and Gotsman, 2021). The integration of atomic transactions in programming languages has also been explored (Borgström et al., 2009)). In terms of formal models of concurrency, the extension of CCS with atomic transactions has been investigated in the past (Acciai et al., 2007; de Vries et al., 2010; De Vries et al., 2010), but without follow-on research, so it seems. While transition systems have been the bedrock of abstract models of computation since the Turing machine, we are not aware of previous attempts to explore atomic transactions within their context.
Previous work on formal implementations of grassroots platforms employed unary transition systems (Shapiro, 2021, 2023b, 2024; Lewis-Pye et al., 2023). While this new definition of a grassroots protocol tries to capture the same intuition as the original one (Shapiro, 2023a), the new definition is crisper. It is also more restrictive in two senses, and more lax in a third:
  1. (1)
    It is specified in terms of configurations and transitions, not runs as in the original definition, so its restriction applies to all configurations, not only to configurations reachable via a run.
  2. (2)

### v3

@@ 4.3 The Three Platforms are Grassroots as Specified
Proof 4.16.
Consider $P\subset P^{\prime}$ and a configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$. This means that all friendships of members of $P$ in $c$ are with other members in $P$. Hence the transaction in $R(P^{\prime})$ in which a member of $P$ establishes friendship with a member of $P^{\prime}\setminus P$ satisfies the definition of interactivity.
@@
Proof 4.18.
Consider $P\subset P^{\prime}$ and a configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$. This means that all coins held by members of $P$ in $c$ are $P$-coins. Hence the transaction in $R(P^{\prime})$ in which a member of $P$ swaps coins with a member of $P^{\prime}\setminus P$ (possibly preceded by transactions in which the two members mint coins, if they have not done so already) satisfies the definition of interactivity.
@@
Proof 4.20.
Consider $P\subset P^{\prime}$ and a configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$. This means that the subgraph of $G$ projected by members of $P$ is a connected component of $G$. Hence the Join transaction in $R(P^{\prime})$ in which a member of $P$ forms an edge with a member of $P^{\prime}\setminus P$, the direction of which depends on the order $\succ$ of their identifiers, satisfies the definition of interactivity.
5 Implementation
The notion of implementations among distributed transition systems, as well as their fault-tolerance, has been studied extensively [1, 27, 34, 32, 42, 50]. Here, we discuss this notion briefly and informally, and plan follow-on work to do so formally.
@@ 5 Implementation
A challenge in this implementation is that a faulty $p$ may fail to either commit or abort following an accept by $q$, leaving $q$ in limbo, at least in regards to this transaction. Solutions to this are a subject of future work.
@@
$k$-ary transactions. The $k$-ary transaction, in which two communities joined or one community leaves another, calls for a consensus protocol executed by the members of the community. To a first approximation, any permissioned consensus (Byzantine Atomic Broadcast) protocol would do. However, the transactions entail changes in the set of agents that participate in consensual decisions, known as reconfiguration in the consensus literature [20, 5, 46]. More generally, the grassroots platform entails multiple dynamically-changing, partially-overlapping, sets of agents engaged in running partially-dependent consensus protocols. This challenge seems to have not been previously addressed in a principled way.
6 Related and Future Work
Atomic transactions have been investigated early in distributed computing, mostly in the context of database systems [28, 31, 30]. Most research since and until today focuses on their efficient and robust implementation [9, 14]. The integration of atomic transactions in programming languages has also been explored [8]). In terms of formal models of concurrency, the extension of CCS with atomic transactions has been investigated in the past [4, 18, 19], but without follow-on research, so it seems. While transition systems have been the bedrock of abstract models of computation since the Turing machine, we are not aware of previous attempts to explore atomic transactions within their context.
Previous work on formal implementations of grassroots platforms employed unary transition systems [42, 44, 45, 29]. While this new definition of a grassroots protocol tries to capture the same intuition as the original one [43], the new definition is crisper. It is also more restrictive in two senses, and more lax in a third:
  1. 1.
    It is specified in terms of configurations and transitions not runs, so its restriction applies to all configurations, not only those reachable via a run.
  2. 2.